# ActAlign: Zero-Shot Fine-Grained Video Classification via Language-Guided Sequence Alignment

### v2

@@ Abstract
We address the task of zero-shot video classification for extremely fine-grained actions (e.g., Windmill Dunk in basketball), where no video examples or temporal annotations are available for unseen classes. While image–language models (e.g., CLIP, SigLIP) show strong open-set recognition, they lack temporal modeling needed for video understanding. We propose ActAlign, a truly zero-shot, training-free method that formulates video classification as a sequence alignment problem, preserving the generalization strength of pretrained image–language models. For each class, a large language model (LLM) generates an ordered sequence of sub-actions, which we align with video frames using Dynamic Time Warping (DTW) in a shared embedding space. Without any video–text supervision or fine-tuning, ActAlign achieves 30.5% accuracy on ActionAtlas—the most diverse benchmark of fine-grained actions across multiple sports—where human performance is only 61.6%. ActAlign outperforms billion-parameter video–language models while using $\sim$8$\times$ fewer parameters. Our approach is model-agnostic and domain-general, demonstrating that structured language priors combined with classical alignment methods can unlock the open-set recognition potential of image–language models for fine-grained video understanding.
Codehttps://github.com/amir-aghdam/ActAlign
@@ 1 Introduction
Understanding fine-grained human activities in video—such as distinguishing a hook shot from a layup in basketball, or recognizing tactical formations in football—requires parsing subtle, temporally ordered visual cues across frames. These fine-grained actions unfold in structured sequences of sub-actions and are often nearly indistinguishable from one another in appearance. In contrast to general activities like swimming, which can often be inferred from a single frame showing a person in water, fine-grained recognition demands attention to temporally extended object interactions, spatial relations, and high-level intent. As such, models must not only understand what is present in a video but also when and how key sub-actions occur. This requires accurately aligning the temporal progression of sub-actions with each fine-grained action to ensure a correct prediction, as shown in Figure 1. While, their high-level label is barely recognized by action recognition models (e.g. such as Roundhouse kick in Kickboxing).
@@
At the same time, contrastive vision–language models such as CLIP (Radford et al. 2021) and SigLIP (Zhai et al. 2023) have demonstrated impressive open-set fine-grained recognition in image domains by training on massive image–text pairs. These contrastive models enable zero-shot classification using natural language prompts and have been widely adopted for downstream recognition tasks. However, extending these capabilities to video understanding introduces new challenges and requires temporal modeling. Existing methods that adapt CLIP-style models to video recognition either average frame-level features (Rasheed et al. 2023; Zohra et al. 2025)—ignoring temporal structure—or fine-tune on target datasets (Wang, Xing, and Liu 2022; Ni et al. 2022; Kim et al. 2024; Wang et al. 2024a), sacrificing generalization and open-set recognition. In both cases, the fine-grained temporal semantics of actions are lost or diluted.
@@
Recent video–language architectures and instruction-tuned LLM-based systems such as Video-LLaMA (Zhang, Li, and Bing 2023), VideoChat (Li et al. 2023) mPLUG-Owl (Ye et al. 2023), Qwen2-VL (Wang et al. 2024b), and DeepSeek-JanusPro (Chen et al. 2025) enable open-ended, dialog-style video understanding through heavy instruction tuning, but they are not tailored for fine-grained video recognition.
@@
Meanwhile, textual grounding (image–text alignment) remains a central challenge in interactive video-language models, especially for open-set and fine-grained video recognition. Dynamic Time Warping (DTW) (Vintsyuk 1968), a classical algorithm for aligning temporally mismatched sequences, has seen renewed interest through differentiable variants (Dogan et al. 2018; Chang et al. 2019) designed for supervised image–text temporal alignment. Yet, these methods rely either on annotated transcripts or example support videos or originally proposed for supervised training, making them impractical for zero-shot recognition. Likewise, approaches using part-level or attribute-level supervision (Wu et al. 2023; Zhu et al. 2024) offer fine-grained cues but lack the ability to model the temporal structure between language-defined actions and visual content.
@@
In this work, we introduce ActAlign, a novel framework that brings the open-set generalization power of image–text models to video classification of extremely fine-grained actions through language-guided sub-action alignment in a truly zero-shot setting. Rather than tuning a model for a specific domain or collapsing the video into a static representation, ActAlign operates in a training-free setting: for each unseen action class, we use a large language model (LLM) to generate a structured sequence of temporal sub-actions that semantically define the class. Then, using the pretrained SigLIP model (Zhai et al. 2023) to extract frame-wise visual and subaction features, we align the frame sequence with the LLM-generated sub-action script via Dynamic Time Warping (DTW) (see Figure 1). This allows us to compute a soft alignment score between different action classes that respects both content and temporal ordering, enabling fine-grained classification in a truly zero-shot manner.
@@ 2.1 Action Recognition
Action recognition has been extensively studied, with early work leveraging hand-crafted features and classic classifiers (Dalal and Triggs 2005; Jain, Jégou, and Bouthemy 2013; Wang and Schmid 2013). The field has since evolved to deep learning approaches (Carreira and Zisserman 2017; Simonyan and Zisserman 2014; Tran et al. 2018; Wang et al. 2016), including CNNs, RNNs, and more recently, Transformer-based models (Arnab et al. 2021; Bertasius, Wang, and Torresani 2021; Kim et al. 2025; Yoshida et al. 2025), which model spatio-temporal dynamics more effectively. Fine-grained action recognition remains challenging due to subtle inter-class variations over time. Recent methods tackle this under supervised (Leong et al. 2022; Sun et al. 2022), semi-supervised (Li, He, and Xu 2022; Dave, Rizve, and Shah 2025), and few-shot (Tang et al. 2023; Hong et al. 2021; Wang et al. 2021a) settings to mitigate annotation costs. Yet, they remain reliant on some level of fine-grained labeling and are constrained to specific domains.
@@ 2.2 Vision-Language Models
Image–Language Models
Foundational image–language models such as CLIP (Radford et al. 2021), SigLip (Zhai et al. 2023), and ALIGN (Jia et al. 2021) learn joint image–text embeddings from large-scale image–caption pairs. Such modelling enables strong open-set recognition without task-specific supervision. These models are widely adopted as pretrained backbones for downstream tasks, including visual question answering (Li, Baldridge, and Hoi 2021; Tsimpoukelli et al. 2021; Li et al. 2022b; Steiner et al. 2024), image captioning and generation (Mokady, Hertz, and Bermano 2021; Wang et al. 2021b; Mahjourian and Nguyen 2025), and few-/zero-shot classification (Zhou et al. 2021; Khattak et al. 2025). However, they lack temporal modeling, which limits their open-set recognition capability to video inputs.
Video–Language Models
Early efforts in video modeling focused on self-supervised pretraining. Models like ActBERT(Zhu and Yang 2020) and VideoBERT(Sun et al. 2019) applied masked language modeling to video frames. This enabled better transfer to video–language tasks through finetuning on paired video–text data (Miech et al. 2020; Lei et al. 2021; Xu et al. 2021; Feichtenhofer, Fan, and He 2022; Li et al. 2022a). Recent progress in LLM reasoning has driven their integration with visual encoders, enabling open-ended video understanding via user prompts. These models (Ye et al. 2023; Li et al. 2023; Zhang et al. 2024; Wang et al. 2024b), such as Video-LLaMA (Zhang, Li, and Bing 2023), combine pretrained visual backbones with chat-centric LLMs to generate spatio-temporal reasoning in conversational settings. While these models excel at open-ended question answering, they require extensive instruction tuning and are not optimized for fine-grained video recognition.
@@ 2.3 Sequence Alignment
Sequence alignment is a long-standing problem in fields such as biology and signal processing. A widely used technique for aligning temporally misaligned signals is Dynamic Time Warping (DTW)(Berndt and Clifford 1994; Keogh and Pazzani 2000; Rath and Manmatha 2003), which computes the optimal warping path between two ordered sequences. Early work extended DTW to video tasks(Huang et al. 2016; Richard, Kuehne, and Gall 2018). For instance, a seminal approach (Bojanowski et al. 2015) proposed aligning videos to ordered scripts by enforcing the temporal order of events. With the rise of deep learning, DTW has been adapted into differentiable forms (Dogan et al. 2018; Chang et al. 2019), enabling greater flexibility and improved learning capacity. A notable example is OTAM (Cao et al. 2020), which leverages DTW for representation learning in few-shot video classification by aligning each video to its support set. However, these methods often assume access to ground-truth transcripts for each video or rely on task-specific supervision, such as paired video–text exemplars.
@@ 3.2 Our Method
Figure 2: Our ActAlign Method Overview. (1) Subaction Generation: Given fine-grained actions (e.g. Basketball Tactics), we prompt an LLM to decompose each action (e.g. Hookshot, JumpShot, Dunk) into a temporal sequence of sub-actions. (2) Temporal Alignment: Video frames are encoded by a frozen pretrained vision encoder and smoothed via a moving‐average filter. Simultaneously, each subaction is encoded by the text encoder. We compute a cosine‐similarity matrix between frame and subaction embeddings, then apply Dynamic Time Warping (DTW) to find the optimal alignment path and normalized alignment score. (3) Class Prediction: We repeat this process for each candidate action m, compare normalized alignment scores $\hat{\gamma}_{\text{video},m}$, and select the action sequence with the highest score as the final prediction.
@@ 3.3 Preliminary Sub-action Generation by LLM
    Return the list $[s_{m,1},\dots,s_{m,K_{m}}]$ in a consistent, structured format.
@@
We enrich each sub-action by embedding it into a contextualized template: This is a video of doing <action name> in <sport domain> with <subaction>. This substantially improves semantic clarity and further disambiguate sub-actions.
@@ 3.4 Visual and Semantic Feature Encoding
Let $\phi_{t}:\mathcal{T}\to\mathbb{R}^{d}$ be the text encoder. For each class $c_{m}$ and its subaction sequence $S_{m}$, we embed each step:
@@ 3.5 Dynamic Time Warping
Real-world footage often contains abrupt scene changes or irrelevant frames (e.g., replays, advertisements) that introduce noise into $Z_{i}$. To mitigate this, we apply a 1D moving-average filter of width $w$ across the temporal dimension:
@@
with boundary conditions handled via zero padding. The kernel width $w$ controls the trade-off between noise reduction and temporal resolution.
@@
[…]
@@
where $\alpha,\beta$ are learned scaling parameters. The resulting $\hat{A}^{(i,m)}$ is used as the input affinity matrix for DTW alignment.
@@
s.t. $\textit{W is a valid warping path between }[1,K_{m}]\textit{ and }[1,T_{i}].$ (10)
@@
$\gamma_{i,m}=\sum_{(k,t)\in W^{(m,i)}}\hat{A}^{(i,m)}_{k,t},$ (12)
$\hat{A}^{(i,m)}$ is the affinity matrix introduced in Eq. 9. To mitigate the bias toward longer warping paths (which can accumulate higher raw scores), we normalize $\gamma_{i,m}$ by the path length, resulting in the average alignment score:
$\hat{\gamma}_{i,m}=\frac{1}{|W^{(i,m)}|}\gamma_{i,m},$ (13)
where $|W^{(i,m)}|$ is the number of matched frame–subaction pairs. Since the similarity values in $\hat{A}^{(i,m)}$ lie in [0,1][0,1][ 0 , 1 ] (due to the sigmoid in Eq. 9), the normalized alignment score $\hat{\gamma}_{i,m}$ also lies in the range [0,1][0,1][ 0 , 1 ].
@@
mPLUG-Owl-Video (Ye et al. 2023) $7B$ 19.4919.4919.49 - -
VideoChat2 (Li et al. 2023) $7B$ 21.2721.2721.27 - -
VideoLLaMA (Zhang, Li, and Bing 2023) $8B$ 22.7122.7122.71 - -
LLaVA-Next-Video (Zhang et al. 2024) $7B$ 22.9022.9022.90 - -
Qwen2–VL (Wang et al. 2024b) $7B$ 30.2430.2430.24 - -
X-CLIP-L/14-16F (Ni et al. 2022) $0.6B$ 16.2616.2616.26 33.7433.7433.74 49.8949.8949.89
SigLIP–so400m (Zhai et al. 2023) (mean-pool) $0.9B$ 22.9422.9422.94 42.2042.2042.20 63.7063.7063.70
+ DTW Alignment (Ours) $0.9B$ 25.7225.7225.72 45.9945.9945.99 66.2666.2666.26
ActAlign (Ours) $0.9B$ $30.29$ $53.01$ $70.27$
@@ 4.1 Experimental Setup
Our approach is dataset- and domain-agnostic, as it relies solely on language-driven sub-action generation from LLMs and cross-modal alignment without task-specific supervision. To rigorously evaluate its generality, we adopt ActionAtlas (Salehi et al. 2024)—to our knowledge, the most diverse and challenging benchmark for fine-grained action recognition across various domains. For each video $V_{i}$, we retain its multiple-choice candidate set $\{c_{i,1},\dots,c_{i,M_{i}}\}$, and replace each class label with an LLM-generated sub-action sequence. (Dataset statistics is provided in Appendix.)
@@
[…]

### v3

@@ Abstract
[…]
Code Link: https://amir-aghdam.github.io/act-align/
@@ 1 Introduction
Understanding fine-grained human activities in video—such as distinguishing a hook shot from a layup in basketball, or recognizing tactical formations in football—requires parsing subtle, temporally ordered visual cues across frames. These fine-grained actions unfold in structured sequences of sub-actions and are often nearly indistinguishable from one another in appearance. In contrast to general activities like swimming, which can often be inferred from a single frame showing a person in water, fine-grained recognition demands attention to temporally extended object interactions, spatial relations, and high-level intent. As such, models must not only understand what is present in a video but also when and how key sub-actions occur. This requires accurately aligning the temporal progression of sub-actions with each fine-grained action to ensure a correct prediction, as shown in Figure 1. While their high-level label is barely recognized by action recognition models (e.g. such as Roundhouse kick in Kickboxing).
@@
At the same time, contrastive vision–language models such as CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023) have demonstrated impressive open-set fine-grained recognition in image domains by training on massive image–text pairs. These models learn a shared image–text latent space, which allows both visual inputs and natural language descriptions to be embedded for direct comparison without task-specific supervision. This enables zero-shot classification using natural language prompts and has been widely adopted for downstream recognition tasks. However, extending these capabilities to video understanding introduces new challenges and requires temporal modeling. Existing methods that adapt CLIP-style models to video recognition either average frame-level features Rasheed et al. (2023); Zohra et al. (2025)—ignoring temporal structure—or fine-tune on target datasets Wang et al. (2022); Ni et al. (2022); Kim et al. (2024); Wang et al. (2024a), sacrificing generalization and open-set recognition. In both cases, the fine-grained temporal semantics of actions are lost or diluted.
@@
Recent video–language architectures and instruction-tuned LLM-based systems such as Video-LLaMA Zhang et al. (2023), VideoChat Li et al. (2023) mPLUG-Owl Ye et al. (2023), Qwen2-VL Wang et al. (2024b), and DeepSeek-JanusPro Chen et al. (2025) enable open-ended, dialog-style video understanding through heavy instruction tuning, but they are not tailored for fine-grained video recognition.
@@
Meanwhile, textual grounding (image–text alignment) remains a central challenge in interactive video-language models, especially for open-set and fine-grained video recognition. Dynamic Time Warping (DTW) Vintsyuk (1968), a classical algorithm for aligning temporally mismatched sequences, has seen renewed interest through differentiable variants Dogan et al. (2018); Chang et al. (2019) designed for supervised image–text temporal alignment. Yet, these methods rely either on annotated transcripts or example support videos or originally proposed for supervised training, making them impractical for zero-shot recognition. Likewise, approaches using part-level or attribute-level supervision Wu et al. (2023); Zhu et al. (2024) offer fine-grained cues but lack the ability to model the temporal structure between language-defined actions and visual content.
@@
In this work, we introduce ActAlign, a novel framework that brings the open-set generalization power of image–text models to video classification of extremely fine-grained actions through language-guided sub-action alignment in a truly zero-shot setting. Rather than tuning a model for a specific domain or collapsing the video into a static representation, ActAlign operates in a training-free setting: for each unseen action class, we use a large language model (LLM) to generate a structured sequence of temporal sub-actions that semantically define the class. Then, using the pretrained SigLIP model Zhai et al. (2023) to extract frame-wise visual and sub-action features, we align the frame sequence with the LLM-generated sub-action script via Dynamic Time Warping (DTW) (see Figure 1). This allows us to compute a soft alignment score between different action classes that respects both content and temporal ordering, enabling fine-grained classification in a truly zero-shot manner.
@@ 2.1 Action Recognition
Action recognition has been extensively studied, with early work leveraging hand-crafted features and classic classifiers Dalal & Triggs (2005); Jain et al. (2013); Wang & Schmid (2013). The field has since evolved to deep learning approaches Carreira & Zisserman (2017); Simonyan & Zisserman (2014); Tran et al. (2018); Wang et al. (2016), including CNNs, RNNs, and more recently, Transformer-based models Arnab et al. (2021); Bertasius et al. (2021); Kim et al. (2025); Yoshida et al. (2025), which model spatio-temporal dynamics more effectively. Fine-grained action recognition remains challenging due to subtle inter-class variations over time. Recent methods tackle this under supervised Leong et al. (2022); Sun et al. (2022), semi-supervised Li et al. (2022c); Dave et al. (2025), and few-shot Tang et al. (2023); Hong et al. (2021); Wang et al. (2021a) settings to mitigate annotation costs. Yet, they remain reliant on some level of fine-grained labeling and are constrained to specific domains.
@@ 2.2 Vision-Language Models
2.2.1 Image–Language Models
Foundational image–language models such as CLIP Radford et al. (2021), SigLIP Zhai et al. (2023), and ALIGN Jia et al. (2021) learn joint image–text embeddings from large-scale image–caption pairs. Such modeling enables strong open-set recognition without task-specific supervision. These models are widely adopted as pretrained backbones for downstream tasks, including visual question answering Li et al. (2021); Tsimpoukelli et al. (2021); Li et al. (2022b); Steiner et al. (2024), image captioning and generation Mokady et al. (2021); Wang et al. (2021b), and few-/zero-shot classification Zhou et al. (2021); Khattak et al. (2025). However, they lack temporal modeling, which limits their open-set recognition capability to video inputs.
2.2.2 Video–Language Models
Early efforts in video modeling focused on self-supervised pretraining. Models like ActBERT Zhu & Yang (2020) and VideoBERT Sun et al. (2019) applied masked language modeling to video frames. This enabled better transfer to video–language tasks through finetuning on paired video–text data Miech et al. (2020); Lei et al. (2021); Xu et al. (2021); Feichtenhofer et al. (2022); Li et al. (2022a). Recent progress in LLM reasoning has driven their integration with visual encoders, enabling open-ended video understanding via user prompts. These models Ye et al. (2023); Li et al. (2023); Zhang et al. (2024); Wang et al. (2024b), such as Video-LLaMA Zhang et al. (2023), combine pretrained visual backbones with chat-centric LLMs to generate spatio-temporal reasoning in conversational settings. While these models excel at open-ended question answering, they require extensive instruction tuning and are not optimized for fine-grained video recognition.
@@ 2.3 Sequence Alignment
Sequence alignment is a long-standing problem in fields such as biology and signal processing. A widely used technique for aligning temporally misaligned signals is Dynamic Time Warping (DTW)Berndt & Clifford (1994); Keogh & Pazzani (2000); Rath & Manmatha (2003), which computes the optimal warping path between two ordered sequences. Early work extended DTW to video tasksHuang et al. (2016); Richard et al. (2018). For instance, a seminal approach Bojanowski et al. (2015) proposed aligning videos to ordered scripts by enforcing the temporal order of events. With the rise of deep learning, DTW has been adapted into differentiable forms Dogan et al. (2018); Chang et al. (2019), enabling greater flexibility and improved learning capacity. A notable example is OTAM Cao et al. (2020), which leverages DTW for representation learning in few-shot video classification by aligning each video to its support set. However, these methods often assume access to ground-truth transcripts for each video or rely on task-specific supervision, such as paired video–text exemplars.
@@ 3.2 Our Method
Figure 2: Our ActAlign Method Overview. (1) Sub-action Generation: Given fine-grained actions (e.g. Basketball Tactics), we prompt an LLM to decompose each action (e.g. Hookshot, JumpShot, Dunk) into a temporal sequence of sub-actions. (2) Temporal Alignment: Video frames are encoded by a frozen pretrained vision encoder and smoothed via a moving‐average filter. Simultaneously, each sub-action is encoded by the text encoder. We compute a cosine‐similarity matrix between frame and sub-action embeddings, then apply Dynamic Time Warping (DTW) to find the optimal alignment path and normalized alignment score. (3) Class Prediction: We repeat this process for each candidate action m, compare normalized alignment scores $\hat{\gamma}_{\text{video},m}$, and select the action sequence with the highest score as the final prediction.
@@ 3.3 Preliminary Sub-action Generation by LLM
    Return the list $s_{m,1},\dots,s_{m,K_{m}}]$ in a consistent, structured format.
@@
We enrich each sub-action by embedding it into a contextualized template: This is a video of doing <action name> in <sport domain> with <sub-action>. This substantially improves semantic clarity and further disambiguates sub-actions. The sport domain is automatically provided in the ActionAtlas metadata (and could also be inferred from the action name itself).
@@ 3.4 Visual and Semantic Feature Encoding
Let $\phi_{t}:\mathcal{T}\to\mathbb{R}^{d}$ be the text encoder. For each class $c_{m}$ and its sub-action sequence $S_{m}$, we embed each step:
@@ 3.5 Dynamic Time Warping
Real-world footage often contains abrupt scene changes or irrelevant frames (e.g., replays, advertisements) that introduce noise into $Z_{i}$. To mitigate this, we apply a simple, parameter-free 1D moving-average filter of width $w$ across the temporal dimension:
@@
with boundary conditions handled via zero padding. The kernel width $w$ controls the trade-off between noise reduction and temporal resolution. The smoothing effectively aggregates short-term temporal context, dampening isolated spikes while preserving the overall action dynamics. As it introduces no learnable parameters or additional training, it remains lightweight and fully compatible with the training-free setting.
@@
where each $\langle\cdot,\cdot\rangle$ is the inner product of $\mathcal{L}2$-normalized vectors, yielding values in [1,1][-1,1]. Following the SigLIP prediction approach, we then apply a sigmoid function $\sigma(\cdot)$ to transform these values into affinity scores in $[0,1]$:
@@
where $\alpha,\beta$ are pre-learned scaling parameters as part of SigLIP’s pretraining recipe. The resulting $\hat{A}^{(i,m)}$ is used as the input affinity matrix for DTW alignment.
@@
[…]
@@
$\gamma_{i,m}=\sum_{(k,t)\in W^{(m,i)}}\hat{A}^{(m,i)}_{k,t}$, (12)
$\hat{A}^{(m,i)}$ is the affinity matrix introduced in Eq. 9. To mitigate the bias toward longer warping paths (which can accumulate higher raw scores), we normalize $\gamma_{i,m}$ by the path length, resulting in the average alignment score:
$\hat{\gamma}_{i,m}=\frac{1}{|W^{(m,i)}|}\gamma_{i,m},$ (13)
where $|W^{(m,i)}|$ is the number of matched frame–sub-action pairs. Since the similarity values in $\hat{A}^{(i,m)}$ lie in $[0,1]$ (due to the sigmoid in Eq. 9), the normalized alignment score $\hat{\gamma}_{i,m}$ also lies in the range $[0,1]$.
@@
mPLUG-Owl-Video Ye et al. (2023) $7B$ $19.49$ - -
VideoChat2 Li et al. (2023) $7B$ $21.27$ - -
VideoLLaMA Zhang et al. (2023) $8B$ $22.71$ - -
LLaVA-Next-Video Zhang et al. (2024) $7B$ $22.90$ - -
Qwen2–VL Wang et al. (2024b) $7B$ $30.24$ - -
X-CLIP-L/14-16F Ni et al. (2022) $0.6B$ $16.26$ $33.74$ $49.89$
SigLIP–so400m Zhai et al. (2023) (mean-pool) $0.9B$ $22.94$ $42.20$ $63.70$
+ DTW Alignment (Ours) $0.9B$ $25.72$ $45.99$ $66.26$
ActAlign (Ours) $0.9B$ $30.40\pm 0.11$ $53.01$ $70.27$
@@ 4.1 Experimental Setup
Our approach is dataset- and domain-agnostic, as it relies solely on language-driven sub-action generation from LLMs and cross-modal alignment without task-specific supervision. To rigorously evaluate its generality, we adopt ActionAtlas Salehi et al. (2024)—to our knowledge, the most diverse and challenging benchmark for fine-grained action recognition across various domains. For each video $V_{i}$, we retain its multiple-choice candidate set $\{c_{i,1},\dots,c_{i,M_{i}}\}$, and replace each class label with an LLM-generated sub-action sequence. (Dataset statistics is provided in the Appendix.)
@@
Following Chen & Huang (2021); Bosetti et al. (2025), we report Top-$k$ accuracy for $k\in\{1,2,3\}$:
$\text{Top-}k=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\Bigl(\text{rank}_{i}(\hat{y}_{i})\leq k\Bigr),$